OpenStudy (anonymous):

which inequality is equivalent to 7x-8<3x?..x<2..x>-2..x<-0.5 or..x>-0.8

OpenStudy (anonymous):

Do you know how to solve this inequality?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

can you solve 2x-8=3?

OpenStudy (anonymous):

oh i mean 7x-8=3

OpenStudy (anonymous):

im a little unsure

OpenStudy (anonymous):

ok first you have to get your varible on one side and your constants on the other right? so we end up with 7x=3+8 => 7x=14

OpenStudy (anonymous):

now divide both sides by 7 and what do you get?

OpenStudy (anonymous):

x<2?

OpenStudy (anonymous):

very correct

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

No prblem just remeber that you look at it like a normal equation and just reinsert the inequality, unless you have to divide by a negitive number then it will switch sides

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

tag